the length of a rectangle is twice its width the area of the rectangle is 648 what are the dimensions (length and width) of the rectangle? what is the perimeter?
Answer by checkley77(12844)

L=2W
AREA=LW
648=LW
648=2W*W
648=2W^2
W^2=648/2
W^2=324
W=SQRT324
W=18 ANS. FOR THE WIDTH.
L=2*18=36 ANS. FOR THE LENGTH.
PROOF:
18*36=648
648=648
P=2L+2W
P=2*18+2*36
P=36+72
P=108 ANS. FOR THE PERIMETER.